A mixture containing only CS2(g) and excess O2(g) at a total pressure of 90 kPa is placed in a sealed vessel. A spark causes the CS2 to ignite, burning it completely, according to the equation: CS2(g)+3O2(g)→CO2(g)+2SO2(g) After the reaction is completed and the vessel is cooled to the initial temperature, the total pressure in the vessel drops to 80 kPa. What was the mole fraction of O2(g) in the initial mixture?


1
Expert's answer
2021-03-12T05:56:35-0500

Here, we can liken the pressure to mole

We would notice that the pressure for oxygen = (90-80)kPa = 10kPa

We can say the fraction of oxygen = 10/90 = 1:9